### v1

# Rethinking LLM-Based Recommendations: A Query Generation-Based, Training-Free Approach

Abstract
Existing large language model (LLM)-based recommendation methods face several challenges, including inefficiency in handling large candidate pools, sensitivity to item order within prompts (“lost in the middle” phenomenon), poor scalability, and unrealistic evaluation with random negative sampling. To effectively address these issues, we propose a Query-to-Recommendation approach, which leverages LLMs to generate personalized queries for retrieving relevant items from the entire candidate pool, eliminating the need for candidate pre-selection. This method can be integrated into an ID-based recommendation system without additional training, enhances recommendation performance and diversity through LLM’s world knowledge, and works well even for minor item groups. Experiments on three datasets show up to 57% performance improvement, with an average gain of 31%, demonstrating competitive zero-shot performance and further gains when ensembled with existing models.
Rethinking LLM-Based Recommendations:
A Query Generation-Based, Training-Free Approach
Donghee Han       Hwanjun Song       Mun Yong Yi KAIST, Republic of Korea {handonghee, songhwanjun, munyi}@kaist.ac.kr Corresponding author
1 Introduction
Recommendation systems play a crucial role in delivering personalized content and service across various domains, including e-commerce, social media, and streaming services. As the demand for more accurate, diverse and user-centric recommendations continues to increase, the integration of large language models (LLM) has emerged as a promising advancement Wu et al. (2024); Zhao et al. (2024). LLMs possess extensive knowledge of the world and exhibit remarkable abilities in understanding and generating text Yang et al. (2024), enabling new opportunities to improve recommendation systems beyond the limitations of traditional collaborative filtering (CF) and content-based methods. Recently, numerous studies have been proposed to leverage LLM in recommendation systems Ramos et al. (2024); Zhang et al. (2025); Lu et al. (2024).
However, existing LLM-based recommendation methods face significant challenges when applied to real-world environments. Most of the methods typically rely on incorporating candidate items directly into the prompt Hou et al. (2024); Kim et al. (2024); Bao et al. (2023); Liao et al. (2024), leading to inefficiencies, especially with large candidate pools that exceed 10K items. These approaches require a separate recommendation model for the selection of candidates from large-scale item pools. Additionally, the "lost in the middle" phenomenon occurs, where the position of items within the prompt affects the recommendation results Hou et al. (2024). Previous studies evaluated on naive random negative sampling do not accurately reflect real-world scenarios Hou et al. (2024); Bao et al. (2023); Kim et al. (2024); Bai et al. (2024). When an additional recommendation model is used for candidate selection, the selected items tend to exhibit higher similarity compared to random samples. It remains unclear whether LLMs can effectively rerank items in such cases.
Figure 1: Comparison between the traditional serialized pipeline and the proposed parallel approach.
Another consideration when leveraging LLMs in recommendation systems is the cost and time required for training. Early study has proposed fine-tuning of LLMs on recommendation datasets to improve the recommendation ability Bao et al. (2023). More recent study used LLM as zero-shot ranker Hou et al. (2024); however, their performance was inferior compared to traditional training-based recommendation models. This result suggest that LLMs struggle to capture CF-signals, leading to the proposal of methods that enhance recommendation performance by learning CF-signals Kim et al. (2024); Dong et al. (2025). In these studies, user and item representations obtained from CF-based recommendation models were used for LLM training, and the training cost was reduced by fixing the LLM and updating only the weights of the additional projection layers. However, these methods still require gradient backpropagation through the LLM, which requires large GPU memory. Moreover, an additional training phase is needed to achieve alignment with the recommendation model.
Through our experiments, we observed that existing reranking-based methods struggle to perform effectively when using other recommendation models as candidate selectors. This indicates a critical need for a new approach that can operate efficiently within large candidate pools and seamlessly adapt to evolving LLMs without fine-tuning. The motivation behind this work lies in the unique strengths of LLMs. Their vast knowledge of the world enables them to provide flexible and context-sensitive recommendations in open-domain and unstructured scenarios Sun et al. (2024).
However, most studies primarily use LLMs as rerankers, which limits their potential by making them dependent on the performance of upstream candidate selectors. As a result, these methods can create bottlenecks due to the misalignment between the CF-based recommendation model and the differing perspectives of LLM. To address this issue, we propose a parallelized approach in which both the LLM and the recommendation model independently extract top-k items from the entire item pool.
Fig.1 illustrates the difference between the traditional serialized approach and our proposed parallel approach. Our proposed parallel architecture maximizes the utilization of the LLM’s world knowledge and creates synergy by effectively merging the two complementary perspectives. We argue that LLMs can offer a complementary perspective to CF-based recommendation models. To fully leverage their potential, LLMs need direct access to the global item pool; however, incorporating the entire item pool into prompts is impractical. To address this, we introduce a query generation mechanism that enables the LLM to formulate feature-aware queries based on user’s history and key features of items.
To address these challenges, we propose Query-to-Recommendation (QueRec) framework, a novel recommendation method that leverages the complementary strengths of LLM and traditional recommendation models. Unlike existing serialized candidate selection LLM-reranking pipelines, our approach uses LLMs to generate personalized queries that retrieve relevant items directly from the global item pool. This zero-shot methodology eliminates the dependency on candidate pre-selection models and additional training, enabling flexible integration of new LLMs as they emerge. By combining CF signals from traditional recommendation models with world knowledge in LLMs, our method fosters a synergistic effect, enhancing both recommendation accuracy and diversity.
In our approach, the LLM generates feature-aware queries based on user history to incorporate its extensive background knowledge into the recommendation process, while the recommendation model contributes CF-based signals. These two components operate in a parallel architecture to enhance recommendation performance. Specifically, the LLM generates item queries from item information to create item embeddings and derives user queries from user history to produce user embeddings. The similarity between these embeddings is used to select top-k items. Finally, these items are reranked by integrating the similarity scores with the top-k items selected by the recommendation model, effectively combining both perspectives for improved recommendations.
Our method not only improves the performance of the recommendation, but also promotes greater diversity and adaptability, effectively addressing the limitations of existing LLM-based recommendation systems. By generating feature-aware queries, it provides a more transparent rationale for item selection than approaches based on user reviews, offering direct insights into why specific items are recommended. The training-free nature of our framework allows for seamless incorporation of new items based on their key features without additional model updates, and it can also handle large, dynamically changing item pools. This advantage is particularly beneficial in real-world e-commerce scenarios, where large volumes of new items are continuously introduced and updated. Consequently, our design significantly reduces both infrastructure costs and maintenance overhead, making the system highly scalable for rapidly evolving recommendation environments.
Our contributions in this work are as follows.
  • We propose a novel and practical Query-to-Recommendation approach that leverages LLMs to generate personalized queries, addressing the limitations of LLM-Reranking methods that rely on candidate sets.
  • Our method achieves competitive zero-shot performance comparable to state-of-the-art recommendation models, without requiring additional training.
  • We demonstrate that integrating our approach with existing methods improves performance, supported by experimental results.
2 Related Work
2.1 Sequential Recommendation Systems
Sequential recommendation has emerged as a prominent paradigm, leveraging users’ recent interactions to predict the next likely item of interest. Deep learning-based models, such as SASRec Kang and McAuley (2018) and BERT4Rec Sun et al. (2019), have demonstrated strong performance in capturing complex sequential user behaviors, effectively modeling temporal dependencies to improve the precision of the recommendation. Recommendation systems that rely solely on item IDs fail to capture the key features of the items. To address this limitation, various sequential recommendation models have been proposed that incorporate additional information Xie et al. (2022); Yuan et al. (2023).
Figure 2: Overview of QueRec:
2.2 LLM-based Recommendation Systems
With the rapid advancement of LLMs, there has been growing interest in incorporating them into recommendation systems Wu et al. (2024). Early LLM-based recommendation approaches fine-tuned pre-trained models such as T5 Raffel et al. (2020) to generate item recommendations by providing item or user IDs and outputting item IDs in a text format Geng et al. (2022); Li et al. (2023). However, this approaches had limitations in fully leveraging the rich knowledge embedded in LLMs. Recently, studies have proposed improving performance by distilling knowledge from larger LLMs Wang et al. (2024b) and incorporating CF-signals from user-item graphs to achieve better performance Wang et al. (2024a).
Another direction in LLM-based recommendation research explores natural language prompts instead of item and user IDs, allowing LLMs to leverage their extensive knowledge and zero-shot capabilities. Early studies fine-tuned LLMs using LoRA Bao et al. (2023), while later approaches froze LLM parameters and updated only additional components to incorporate CF-signals more efficiently Kim et al. (2024); Dong et al. (2025). Recent work introduced zero-shot LLM-based reranking Hou et al. (2024), but these methods remain limited by reliance on upstream candidate selectors, inefficiencies in large candidate pools, and the "lost in the middle" phenomenon. To address these challenges, we propose a method that better aligns with real-world recommendation settings, ensuring efficiency and scalability in large-scale item pools.
3 Method: QueRec Framework
3.1 Overview
We introduce Query-to-Recommendation (QueRec), a novel framework designed to leverage the complementary strengths of LLMs through large pool of items, while effectively connecting the diverse knowledge embedded in LLMs with the features of the elements. The proposed method leverages the world knowledge of LLMs to generate personalized queries, which can be effectively integrated with traditional recommendation models to improve the accuracy and diversity of recommendations. Fig.2 illustrates the overall process of our method. The overall process involves generating queries from item information to create item queries, leveraging user history to generate user queries, retrieving top-k items based on similarity scores between user queries and item queries, and finally applying a reranking mechanism that fuses insights from both the LLM and the recommendation model. This architecture enables robust, diverse, and personalized recommendations without the need for extensive training of LLMs, and direct search is available in a large item pool, making it scalable for real-world applications.
3.2 Item Query Generation
To effectively leverage the world knowledge of LLMs to highlight key features of items and align them with user preferences, we employ LLMs to generate queries that reflect the distinctive features of each item. This process involves constructing prompts using item metadata (e.g. title, category, description) and user reviews to provide contextual information to the LLM. Based on these prompts, the LLM generates 10 diverse queries that capture different aspects of the item, leveraging its extensive knowledge. Finally, the generated queries are combined with the original item metadata to create a comprehensive item representation.
$q_{i}=\text{LLM}(p_{i},m_{i},r_{i}),\quad d_{i}=m_{i}\oplus q_{i}$ (1)
Equation (1) describes the process of extracting item queries and constructing the complete item representation corpus. Given an item $i$, we construct a query $q_{i}$ using the LLM with inputs: item query generation prompt $p_{i}$, item metadata $m_{i}$, and user reviews $r_{i}$. The final textual representation $d_{i}$ is formed by concatenating the metadata of the item with the generated query. Here, $\oplus$ represents the concatenation operator, integrating both structured metadata and insights derived from LLM to improve the retrieval effectiveness of items.
3.3 User Query Generation
The user representation is generated by incorporating information from the user’s historical interactions and preferences. This process begins with the construction of prompts that combine user history (including item titles and metadata), user reviews, and previously generated item queries to provide rich contextual information. Based on these prompts, the LLM generates 10 personalized queries that reflect the user’s interaction history. The generated queries are then refined by incorporating the most recently interacted item information, ensuring that recent user preferences are effectively captured.
$q_{u}=\text{LLM}(p_{u},h_{u},r_{u},q_{l}),\quad d_{u}=m_{l}\oplus q_{u}$ (2)
Equation (2) describes the process of constructing the textual user representation. Given a user $u$, we construct a query $q_{u}$ using the LLM with inputs: user query generation prompt $p_{u}$, user history $h_{u}$ (which includes item titles and metadata), user reviews $r_{u}$, and the most recent item query $q_{l}$. The final user representation $d_{u}$ is obtained by concatenating the user history with the generated query and incorporating the last interacted item metadata $m_{l}$ to better reflect temporal user preferences. The prompt format and examples are described in the Appendix C .
3.4 Similarity-based Retrieval
With the user and item textual representations obtained from the previous steps, we perform similarity-based retrieval:
$v_{u}$ $=\text{Enc}(d_{u}),\quad v_{i}=\text{Enc}(d_{i})$ (3)
$s_{u,i}$ $=\cos(v_{u},v_{i})$
^$\hat{\mathcal{I}}_{u}=\underset{i\in\mathcal{I}}{\text{arg max }s_{u,i}},\quad\text{where }|\hat{\mathcal{I}}_{u}|=k$ (4)
Equation (3) defines the embedding extraction via the pre-trained text encoder Enc and similarity computation between the user embedding $v_{u}$ and item embeddings $v_{i}$. We compute cosine similarity $s_{u,i}$ between the embeddings to measure their relevance. Based on these scores, we select the top-k items ^$\hat{\mathcal{I}}_{u}$ for the user query, as shown in Equation (4), which are the most relevant items to the user’s preferences.
3.5 Divergent Perspective Reranking
To effectively integrate LLM-based semantic insights and CF-based collaborative signals, we propose a Divergent Perspective Reranking method. Initially, similarity scores from both models (LLM and CF) are normalized into the [0,1] range via min-max normalization, ensuring fair comparison across methods:
$\tilde{s}_{u,i}^{X}=\frac{s_{u,i}^{X}-\min(s_{u}^{X})}{\max(s_{u}^{X})-\min(s_{u}^{X})},$ (5)
$X\in\{LLM,CF\}.$
Subsequently, we calculate the final score using a convex combination (CC) of the normalized scores:
$s_{u,i}^{*}=\lambda\tilde{s}_{u,i}^{LLM}+(1-\lambda)\tilde{s}_{u,i}^{CF}.$ (6)
The initial $\lambda$ is determined by the validation Hit@10 performance of each model:
$\lambda_{\text{init}}=\frac{H_{10}^{LLM}}{H_{10}^{LLM}+H_{10}^{CF}}.$ (7)
Further analysis revealed that the intersection ratio ($\omega$) between hit items from QueRec and CF models varies significantly across datasets and models. In the case of low intersection ratio with imbalanced weights (e.g., 0.3) causes interference between the two recommended item lists, resulting in performance decreasing. To address this, we adjust $\lambda$ by incorporating the intersection ratio:
$\lambda=\omega\cdot\lambda_{\text{init}}+(1-\omega)\cdot 0.5.$ (8)
When overlap is minimal ($\omega\approx 0$), $\lambda$ naturally converges to 0.5, reflecting uncertainty and ensuring a balanced influence from both sources. This adjustment enhances recommendation quality by effectively merging diverse perspectives from LLM and CF knowledge.
4 Experiment
4.1 Datasets and Metrics
Dataset #Users #Items #Reviews Density (%)
Sports 35,598 18,357 296,337 0.0453
Beauty 22,363 12,101 198,502 0.0734
Toys 19,412 11,924 167,597 0.0724
Table 1: Statistics of the datasets.
We conducted experiments on three widely used benchmark datasets collected from the Amazon e-commerce platform: Sports & Outdoors, Beauty, and Toys & Games. We selected the three aforementioned domains for experimentation. Each dataset consists of user interactions, including a user ID, an item ID, a rating, a review, and a timestamp. The statistics of the data set are provided in Table 1. Following previous studies Li et al. (2023); Wang et al. (2024a), we partitioned the data in the same way that the last interaction of a user served as the test label, the second to last as the validation label, and the remaining interactions as training data.
To evaluate the performance of the model, we adopted two widely used ranking metrics: Hit Rate (HR) and Normalized Discounted Cumulative Gain (NDCG), with $k\in\{5,10\}$.
Model Sports Beauty Toys
H@5 N@5 H@10 N@10 H@5 N@5 H@10 N@10 H@5 N@5 H@10 N@10
SASRec 0.0328 0.0179 0.0499 0.0234 0.0561 0.0323 0.0872 0.0423 0.0628 0.0354 0.0915 0.0447
+ LLMRank ICL 0.0201 0.0099 0.0341 0.0144 0.0316 0.0156 0.0555 0.0233 0.0410 0.0204 0.0636 0.0277
+ LLMRank Seq 0.0187 0.0088 0.0328 0.0134 0.0279 0.0134 0.0506 0.0208 0.0388 0.0190 0.0624 0.0267
+ LLMRank Recent 0.0198 0.0099 0.0335 0.0143 0.0304 0.0154 0.0525 0.0225 0.0415 0.0213 0.0639 0.0286
+ TALLRec (w/o FT) 0.0092 0.0046 0.0172 0.0072 0.0077 0.0038 0.0145 0.0060 0.0236 0.0122 0.0409 0.0178
+ TALLRec (w FT) 0.0036 0.0028 0.0054 0.0033 0.0048 0.0029 0.0085 0.0041 0.0097 0.0069 0.0125 0.0078
+ A-LLMRec 0.0200 0.0097 0.0435 0.0174 0.0063 0.0029 0.0311 0.0109 0.0493 0.0243 0.0911 0.0380
+ QueRec 0.0367 0.0242 0.0571 0.0307 0.0675 0.0462 0.1022 0.0574 0.0856 0.0594 0.1189 0.0701
Improvement 11.89% 35.20% 14.43% 31.20% 20.32% 43.03% 17.20% 35.70% 36.31% 67.80% 29.95% 56.82%
DIF-SR 0.0360 0.0197 0.0542 0.0256 0.0576 0.0336 0.0901 0.0442 0.0700 0.0404 0.0995 0.0499
+ LLMRank ICL 0.0210 0.0104 0.0365 0.0154 0.0323 0.0159 0.0545 0.0231 0.0393 0.0197 0.0643 0.0278
+ LLMRank Seq 0.0185 0.0088 0.0341 0.0138 0.0289 0.0139 0.0509 0.0210 0.0384 0.0189 0.0626 0.0268
+ LLMRank Recent 0.0204 0.0102 0.0356 0.0151 0.0318 0.0161 0.0532 0.0230 0.0433 0.0223 0.0656 0.0296
+ TALLRec (w/o FT) 0.0168 0.0095 0.0325 0.0145 0.0239 0.0166 0.0422 0.0224 0.0307 0.0173 0.0577 0.0260
+ TALLRec (w FT) 0.0182 0.0150 0.0188 0.0152 0.0182 0.0150 0.0188 0.0152 0.0191 0.0115 0.0316 0.0156
+ A-LLMRec 0.0240 0.0122 0.0449 0.0191 0.0093 0.0046 0.0328 0.0121 0.0362 0.0151 0.0896 0.0328
+ QueRec 0.0371 0.0247 0.0590 0.0318 0.0699 0.0472 0.1032 0.0579 0.0885 0.0615 0.1228 0.0726
Improvement 3.06% 25.38% 8.86% 24.22% 21.35% 40.48% 14.54% 31.00% 26.43% 52.23% 23.42% 45.49%
ELMRec 0.0492 0.0414 0.0569 0.0437 0.0610 0.0503 0.0729 0.0540 0.0706 0.0616 0.0749 0.0623
+ LLMRank ICL 0.0229 0.0128 0.0312 0.0155 0.0290 0.0157 0.0409 0.0196 0.0296 0.0169 0.0367 0.0192
+ LLMRank Seq 0.0154 0.0084 0.0279 0.0124 0.0236 0.0128 0.0355 0.0166 0.0279 0.0162 0.0366 0.0190
+ LLMRank Recent 0.0167 0.0091 0.0272 0.0125 0.0236 0.0129 0.0361 0.0169 0.0275 0.0160 0.0346 0.0183
+ TALLRec (w/o FT) 0.0214 0.0122 0.0277 0.0142 0.0221 0.0128 0.0309 0.0156 0.0290 0.0173 0.0352 0.0194
+ TALLRec (w FT) 0.0151 0.0086 0.0207 0.0104 0.0084 0.0047 0.0112 0.0056 0.0259 0.0155 0.0315 0.0173
+ A-LLMRec 0.0007 0.0003 0.0026 0.0009 0.0000 0.0000 0.0004 0.0001 0.0005 0.0002 0.0012 0.0004
+ QueRec 0.0589 0.0455 0.0746 0.0506 0.0816 0.0625 0.1060 0.0704 0.1092 0.0777 0.1350 0.0860
Improvement 19.72% 9.90% 31.11% 15.79% 33.77% 24.25% 45.40% 30.37% 54.67% 26.14% 80.24% 38.04%
Table 2: Evaluation results in the RecSys-LLM cooperation scenario. The best and second-best results for each metric are highlighted in bold and underlined, respectively. "H@k" and "N@k" denote Hit Rate and NDCG at rank k, respectively.
4.2 Experimental Setup and Baselines
To evaluate the effectiveness of QueRec, we integrated it into an existing CF-based recommender system and compared its performance against existing LLM-based reranking methods. Our goal was to examine whether the proposed parallel architecture would be more effective than the traditional serial reranking approach. For scalability, we employed LLaMA3.2 3B Dubey et al. (2024) for query generation and the implementation of LLM-based baselines, while for similarity-based retrieval, we leveraged a widely used pre-trained encoder Li and Li (2023) without any additional training.
In the first experiment, we evaluated the performance of QueRec in comparison to existing LLM-based reranking methods when applied to a representative sequential recommender system. We selected SASRec Kang and McAuley (2018), DIF-SR Xie et al. (2022), and ELMRec Wang et al. (2024a) as baselines. The existing LLM-based methods perform a reranking of the top 20 items selected by these CF-based models. LLM-based reranking methods include three variations of the LLMRank Hou et al. (2024) without training: sequential, in-context learning, and recency-focused, as well as TALLRec Bao et al. (2023) in two configurations: zero shot and fine-tuned.
In the second experiment, we compared QueRec as a standalone method against traditional training-based approaches. To better reflect real-world recommendation scenarios with a large candidate pool, we conducted all experiments without separate negative item selection, evaluating performance over the entire item pool.
More details on the experimental settings and baselines are provided in Appendix and in the online repository [^1] https://anonymous.4open.science/r/querec/.
4.3 Main Results
Model Sports Beauty Toys
H@5 N@5 H@10 N@10 H@5 N@5 H@10 N@10 H@5 N@5 H@10 N@10
Caser 0.0116 0.0072 0.0194 0.0097 0.0205 0.0131 0.0347 0.0176 0.0166 0.0107 0.0270 0.0141
GRU4Rec 0.0129 0.0086 0.0204 0.0110 0.0164 0.0099 0.0283 0.0137 0.0097 0.0059 0.0176 0.0084
HGN 0.0189 0.0120 0.0313 0.0159 0.0325 0.0206 0.0512 0.0266 0.0321 0.0221 0.0497 0.0277
SASRec 0.0328 0.0179 0.0499 0.0234 0.0561 0.0323 0.0872 0.0423 0.0628 0.0354 0.0915 0.0447
BERT4Rec 0.0115 0.0075 0.0191 0.0099 0.0203 0.0124 0.0347 0.0170 0.0116 0.0071 0.0203 0.0099
FDSA 0.0182 0.0122 0.0288 0.0156 0.0267 0.0163 0.0407 0.0208 0.0228 0.0140 0.0381 0.0189
S3-Rec 0.0251 0.0161 0.0385 0.0204 0.0387 0.0244 0.0647 0.0327 0.0443 0.0294 0.0700 0.0376
DIF-SR 0.0360 0.0197 0.0542 0.0256 0.0576 0.0336 0.0901 0.0442 0.0700 0.0404 0.0995 0.0499
P5 0.0272 0.0169 0.0361 0.0198 0.0503 0.0370 0.0659 0.0421 0.0648 0.0567 0.0709 0.0587
TIGER 0.0264 0.0181 0.0400 0.0225 0.0454 0.0321 0.0648 0.0384 0.0521 0.0371 0.0712 0.0432
QueRec 0.0280 0.0188 0.0403 0.0228 0.0501 0.0354 0.0675 0.0410 0.0680 0.0472 0.0937 0.0555
w/o recent item info 0.0274 0.0185 0.0407 0.0228 0.0459 0.0319 0.0622 0.0372 0.0625 0.0434 0.0908 0.0525
w/o item query 0.0248 0.0165 0.0369 0.0204 0.0454 0.0317 0.0625 0.0373 0.0625 0.0423 0.0888 0.0508
w/o user query 0.0095 0.0061 0.0166 0.0084 0.0211 0.0131 0.0325 0.0168 0.0254 0.0162 0.0412 0.0213
Table 3: Performance comparison of sequential recommendation models. Metrics with lower performance than QueRec are shaded in blue, while higher-performing cases are shaded in red.
The first experiment evaluated the effectiveness of integrating the proposed method into an existing recommendation system to enhance its performance. The results of this experiment are presented in Table 2. We observed that previously proposed LLM-based reranking methods lead to a decrease in performance. This decline seems to be due to the fact that the top 20 candidate items provided by the CF-based model are very similar to each other, which may confuse the LLM during reranking. Additionally, in the case of TALLRec, performance deteriorated in most scenarios when fine-tuning was applied. These results suggest that learning CF-signals solely through natural language-based training remains a challenging task, even when the LLM is trained on candidate items from CF-based models.
In contrast, QueRec achieved an average performance improvement of 31% and a maximum improvement of 57%. In particular, QueRec does not require additional training when integrated into an existing recommendation system, highlighting its efficiency and scalability. These results demonstrate that QueRec can effectively leverage the world knowledge inherent in LLM, thereby providing differentiated recommendations from CF-based models.
(a) Toys.
(b) Beauty.
(c) Sports.
Figure 3: Diversity of recommendations across different datasets and models. The x-axis represents the IDs of the recommended items, while the y-axis indicates the frequency of recommendation. The top-20 recommended items for each user were extracted and compared.
In the second experiment, we compared the performance of QueRec as a standalone method against traditional sequential recommendation models. The results presented in Table 3 demonstrate that QueRec achieves competitive zero shot performance without additional training, outperforming six of the eight trained traditional recommendation models. This finding is particularly notable given that previous studies have indicated that zero-shot LLM-based approaches have failed to surpass full-shot models. These results suggest that QueRec effectively leverages the knowledge and reasoning capabilities of LLMs, providing a significant advantage in generating queries compared to existing reranking methods.
In addition, we conducted an ablation study to assess the contribution of each component in our approach. As shown in Table 3, we evaluated performance changes when each of the following three components was removed: (1) recent item information, (2) item queries, and (3) user queries. Among these, the absence of user queries resulted in the most significant performance degradation, while excluding the other components also led to notable declines. These findings indicate that all proposed components contribute to the improvement in performance, confirming the effectiveness of QueRec.
4.4 In-depth Analysis
In this subsection, we analyze the distinguishing characteristics and advantages of the proposed method. Since QueRec relies solely on pre-trained models without additional training, it effectively leverages the diverse knowledge embedded in LLMs while avoiding bias toward any specific domain.
4.4.1 Diversity of Recommendations
Our proposed method does not rely on domain-specific training, allowing unbiased recommendations, which in turn leads to more diverse item suggestions. Traditional recommendation systems learn not only collaborative signals, but also item popularity, through training. Although this can improve overall performance metrics, it can also lead to unwanted side effects, such as filter bubbles, where frequently recommended items continue to dominate recommendations Nguyen et al. (2014). To evaluate whether the proposed method improves the diversity of recommendations, we compare the distribution of recommended item IDs with those from conventional recommendation systems.
Fig. 3 illustrates the distribution of recommended items between QueRec and the baselines. Both SASRec and DIF-SR exhibit a bias toward lower-ID item groups, suggesting that the models amplify the bias present in the dataset through the training process. In contrast, ELMRec demonstrates distinct distributions, yet still exhibits skewness toward specific items. This bias is likely a result of its multitask learning in a text-to-text format, but it still struggles to maintain recommendation diversity. In contrast, our proposed method, QueRec, shows a more balanced distribution of recommended items, indicating a stronger trend towards diverse recommendations. This characteristic has practical implications for real-world recommendation systems, as it can mitigate filter bubbles and facilitate a more exploratory recommendation experience based on user preferences. These findings highlight that QueRec not only incorporates a distinct perspective compared to conventional recommendation models, but also has the potential to generate synergies when integrated with existing approaches. A more detailed analysis of this can be found in the Appendix B.2.
4.4.2 Performance in Minor Items
Figure 4: Performance comparison on minor item group in three datasets.
To examine whether the observed improvement in the diversity of recommendations improves the performance of minor items, we performed an analysis using the 20% of items that appeared less frequently in the training set as the minor group. Fig. 4 presents the performance of the recommendation of each method for these elements. In most cases, our proposed method outperformed existing approaches, demonstrating its effectiveness in improving recommendations for items observed less frequently.
However, in the Sports dataset, QueRec exhibited performance comparable to that of ELMRec, likely due to ELMRec’s exposure to a broader range of tasks during training. Given that QueRec does not require additional training, these results offer valuable insights into how our method can mitigate the limitations of traditional recommendation systems while leveraging LLMs’ inherent knowledge and generalization capabilities.
5 Conclusion
In this work, we introduced QueRec, a novel framework that enhances the diversity and precision of recommendations by using LLM to generate feature-aware queries. Unlike LLM-based reranking approaches rely on pre-selected candidate sets, QueRec enables direct retrieval from the entire item pool, mitigating performance bottlenecks while achieving competitive zero-shot performance that often exceeding trained baselines. Furthermore, our analysis highlights the ability of the method to reduce bias, improve recommendations for minor items, and ultimately foster a more diverse recommendation. By integrating the complementary strengths of LLMs and recommendation models, QueRec provides a promising direction for future recommendation research, offering a practical, training-free solution adaptable to evolving LLMs and large-scale dynamic item pools.
6 Limitations
Despite its advantages, QueRec has several limitations. First, while it eliminates the need for explicit candidate selection, the effectiveness of query generation depends on the quality of the pre-trained LLM and its ability to generalize across domains. In cases where the LLM lacks sufficient domain-specific knowledge, the generated queries may not capture user intent effectively. Second, since our approach relies on retrieving items based on semantic similarity, it may underperform in domains where structured interaction data (e.g., collaborative filtering signals) are more informative than textual item representations. Finally, while QueRec enhances diversity, it does not explicitly control for fairness, popularity bias, or serendipity, which remain open challenges for future research. Addressing these limitations through hybrid approaches, enhanced query generation mechanisms, and domain-aware enhancements presents a promising direction for further improving LLM-driven recommendation systems.
7 Ethics Statement
Potential Risks
Our study was conducted on fixed datasets and the potential impact of the user in real-world applications has not been examined. Therefore, caution is required when applying the proposed method beyond the controlled experimental setting.
Use of Scientific Artifacts
Our research leveraged open source tools, including PyTorch Paszke et al. (2019), along with pre-trained language models such as LLaMA3.2 and T5 obtained via the Huggingface Wolf et al. (2019) library.
Use of Ai Assistants
We only used ChatGPT to provide a better expression and to refine the language. Some of the code used in the experiment was written with the assistance of Copilot.
Appendix A Baselines
A.1 Recommendation Model
  • Caser Tang and Wang (2018): A CNN-based sequential recommendation model that learns item embeddings through convolution operations to effectively model users’ sequential patterns.
  • GRU4Rec Hidasi (2015): An RNN-based sequential recommendation model that leverages GRU to capture users’ sequential patterns.
  • HGN Ma et al. (2019): A sequential recommendation model that captures both long-term and short-term user interests through a hierarchical gating mechanism, which selectively processes item features and instances while explicitly modeling item-item relationships.
  • SASRec Kang and McAuley (2018): A sequential recommendation model that employs a self-attention mechanism to learn users’ sequential behavior patterns.
  • BERT4Rec Sun et al. (2019): A sequential recommendation model that utilizes BERT architecture to capture users’ sequential interaction patterns.
  • FDSA Zhang et al. (2019): A sequential recommendation model that captures both item-level and feature-level transition patterns by applying separate self-attention blocks to model their relationships, enhancing recommendation performance. It integrates heterogeneous item features using attention mechanisms and combines item and feature transitions to improve next-item prediction.
  • S3-Rec Zhou et al. (2020): A self-attentive model that enhances sequential recommendation by leveraging self-supervised pre-training objectives to capture correlations among attributes, items, subsequences, and sequences using mutual information maximization.
  • DIF-SR Xie et al. (2022): A sequential recommendation model that enhances side information fusion by shifting it from the input to the attention layer, decoupling attention calculations to mitigate rank bottlenecks and improve gradient flexibility, thereby enhancing modeling capacity and boosting recommendation performance.
  • ELMRec Wang et al. (2024a): A T5-based recommendation by enhancing whole-word embeddings to better interpret high-order user-item interactions without graph pre-training, while also addressing recency bias through a reranking mechanism to improve both direct and sequential recommendations.
  • P5 Geng et al. (2022): A T5-base text-to-text recommendation model, converting all recommendation-related data into natural language sequences and utilizing personalized prompts for various recommendation tasks.
  • TIRGER Rajput et al. (2023): A generative model-based recommendation method, explicitly using numeric item-IDs to perform next-item prediction tasks. Unlike our approach, it does not leverage natural language generation, focusing instead on digit-based input-output structures.
To obtain similarity scores for SASRec and DIF-SR, we trained the models using Recbole [^2] https://recbole.io/, a well-validated open-source recommendation framework. Detailed hyperparameters and training results will be made available through our public repository. For ELMRec, we excluded test targets affected by label leakage in the explanation task to ensure fair evaluation. In case of ELMRec, we obtain beam search scores and use them as the recommendation score. The model was trained using the same T5-small backbone and hyperparameters as in prior studies. For all other baselines, we report the performance metrics as documented in previous studies Zhou et al. (2020); Li et al. (2023); Wang et al. (2024a).
A.2 LLM-based Reranking Method
  • LLMRank Hou et al. (2024): A zero-shot reranking approach using LLMs has been proposed to rerank selected candidate items. Three types of prompts were introduced: sequential, in-context learning, and recency-focused.
  • TALLRec Bao et al. (2023): An approach that fine-tunes LLMs using LoRA to enhance their recommendation capabilities, enabling efficient adaptation to recommendation tasks while reducing computational overhead. The initially proposed method used a binary prediction approach, whereas we modified the prompt to select the top-10 items from the candidate set used in a previous study Kim et al. (2024). During training, the prompt included the top-20 candidate items recommended by the traditional recommendation model, while the training target was a reordered list of 10 item titles, ensuring that the label item appeared first. This setup was designed to allow a fair comparison that closely aligns with our proposed method. We trained the LLaMA 3.2 3B with LoRA for 2 epochs with a batch size of 4 and a learning rate of 5e-5.
Appendix B Additional Experiments
In this section, we present additional experiments that were not included in the main text due to space constraints.
B.1 LLM Variation
To evaluate the generalizability of our proposed approach across different LLM architectures, we conducted experiments using various models, including LLaMA 3.1 8B, LLaMA 3.3 70B, Gemma 2 2B, and Gemma 2 3B. Table 4 presents the performance of QueRec in a standalone setting. For the 70B model, we utilized Llama-3.3-70B-Instruct-Turbo in DeepInfra [^3] https://deepinfra.com/ for generation, as it was not feasible to run on our locally available GPUs.
The results demonstrate that our method maintains stable performance across different LLMs, confirming its adaptability to various model architectures. Additionally, we observe that even lightweight models can achieve performance comparable to larger models, highlighting the efficiency and scalability of our approach.
Model Sports Beauty Toys
H@5 N@5 H@10 N@10 H@5 N@5 H@10 N@10 H@5 N@5 H@10 N@10
LLaMA 3.2 3B 0.0280 0.0188 0.0403 0.0228 0.0501 0.0354 0.0675 0.0410 0.0680 0.0472 0.0937 0.0555
LLaMA 3.1 8B 0.0270 0.0185 0.0397 0.0225 0.0491 0.0346 0.0667 0.0402 0.0665 0.0466 0.0926 0.0550
LLaMA 3.3 70B 0.0269 0.0182 0.0399 0.0224 0.0505 0.0361 0.0686 0.0419 0.0678 0.0469 0.0935 0.0552
Gemma 2 2B 0.0272 0.0186 0.0395 0.0225 0.0496 0.0356 0.0682 0.0416 0.0661 0.0457 0.0927 0.0543
Gemma 2 9B 0.0247 0.0165 0.0390 0.0211 0.0481 0.0345 0.0680 0.0409 0.0664 0.0460 0.0940 0.0549
Table 4: Performance variation of QueRec across different LLMs. "H@k" and "N@k" denote Hit Rate and NDCG at rank k, respectively.
To investigate whether previously proposed reranking methods become more effective when utilizing a larger model, we conducted experiments using the LLaMA 3.3 70B. Table 5 presents the results of these experiments. Compared to lightweight models, we observed an overall improvement in performance compare to the cases used LLaMA 3.2 3B, and in some metrics, performance improvements were also observed compared to the standalone recommendation model. However, in most cases, performance degradation was still evident, reaffirming that the reranking approach remains less effective than our proposed method. These findings highlight that simply scaling up the model does not necessarily resolve the limitations of existing reranking strategies.
Model Sports Beauty Toys
H@5 N@5 H@10 N@10 H@5 N@5 H@10 N@10 H@5 N@5 H@10 N@10
DIF-SR 0.0360 0.0197 0.0542 0.0256 0.0576 0.0336 0.0901 0.0442 0.0700 0.0404 0.0995 0.0499
+ LLMRank ICL 0.0315 0.0215 0.0409 0.0246 0.0453 0.0322 0.0501 0.0339 0.0663 0.0478 0.0843 0.0537
+ LLMRank Seq 0.0348 0.0246 0.0431 0.0274 0.0489 0.0361 0.0534 0.0376 0.0725 0.0546 0.0891 0.0601
+ LLMRank Recent 0.0354 0.0252 0.0435 0.0279 0.0491 0.0364 0.0540 0.0380 0.0739 0.0556 0.0904 0.0610
+ QueRec 0.0389 0.0243 0.0604 0.0312 0.0715 0.0479 0.1059 0.0589 0.0868 0.0618 0.1231 0.0735
Table 5: Experimental results on LLaMA 3.3 70B. The best and second-best results for each metric are highlighted in bold and underlined, respectively. "H@k" and "N@k" denote Hit Rate and NDCG at rank k, respectively.
B.2 Item Distribution
We conducted a more detailed analysis of the bias present in conventional recommendation systems. Fig. 5 compares the distribution of item IDs in the training and test datasets used in our experiments. While the overall distributions are similar, we observe that the training dataset is more skewed towards lower item IDs, whereas the test dataset contains a relatively higher frequency of items with larger IDs.
To quantify this observation, we measured the skewness of the distributions. Fig. 6 presents a comparison of the skewness values for both datasets and the recommendation outputs of various models. Our analysis confirms that the training set exhibits a higher degree of skewness compared to the test set. Additionally, traditional recommendation models demonstrate an even greater level of skewness in their recommendation results than the training dataset itself.
These findings indicate that conventional recommendation approaches may contribute to reduced diversity in recommendations. This further underscores the necessity of unbiased recommendation methods, such as our proposed approach, to mitigate such biases and improve recommendation fairness.
Figure 5: Target item ID distribution across different datasets. The x-axis represents the IDs of the target items, while the y-axis indicates the frequency of items.
Figure 6: Comparison of item distribution skewness across datasets and the skewness of recommendation results for each model.
B.3 Divergent Perspective Reranking
Model Sports Beauty Toys
H@5 N@5 H@10 N@10 H@5 N@5 H@10 N@10 H@5 N@5 H@10 N@10
SASRec 0.0328 0.0179 0.0499 0.0234 0.0561 0.0323 0.0872 0.0423 0.0628 0.0354 0.0915 0.0447
+ QueRec CC (Ours) 0.0367 0.0242 0.0571 0.0307 0.0675 0.0462 0.1022 0.0574 0.0856 0.0594 0.1189 0.0701
+ QueRec CC (0.9) 0.0272 0.0179 0.0388 0.0216 0.0488 0.0340 0.0673 0.0400 0.0681 0.0458 0.0949 0.0545
+ QueRec CC (0.7) 0.0317 0.0209 0.0452 0.0253 0.0565 0.0391 0.0816 0.0472 0.0796 0.0538 0.1089 0.0632
+ QueRec CC (0.5) 0.0347 0.0231 0.0534 0.0291 0.0638 0.0434 0.0988 0.0547 0.0837 0.0588 0.1180 0.0699
+ QueRec CC (0.3) 0.0371 0.0219 0.0576 0.0285 0.0655 0.0421 0.0988 0.0529 0.0774 0.0521 0.1107 0.0628
+ QueRec CC (0.1) 0.0345 0.0192 0.0522 0.0249 0.0597 0.0360 0.0907 0.0460 0.0675 0.0401 0.0978 0.0498
+ QueRec RRF 0.0330 0.0198 0.0526 0.0261 0.0585 0.0374 0.0899 0.0475 0.0730 0.0473 0.1056 0.0578
DIF-SR 0.0360 0.0197 0.0542 0.0256 0.0576 0.0336 0.0901 0.0442 0.0700 0.0404 0.0995 0.0499
+ QueRec CC (Ours) 0.0371 0.0247 0.0590 0.0318 0.0699 0.0472 0.1032 0.0579 0.0885 0.0615 0.1228 0.0726
+ QueRec CC (0.9) 0.0274 0.0180 0.0389 0.0217 0.0489 0.0340 0.0676 0.0401 0.0690 0.0461 0.0949 0.0545
+ QueRec CC (0.7) 0.0327 0.0214 0.0465 0.0259 0.0580 0.0402 0.0833 0.0482 0.0804 0.0541 0.1098 0.0635
+ QueRec CC (0.5) 0.0360 0.0239 0.0557 0.0303 0.0668 0.0457 0.0982 0.0558 0.0842 0.0594 0.1204 0.0711
+ QueRec CC (0.3) 0.0380 0.0227 0.0580 0.0292 0.0654 0.0437 0.1003 0.0550 0.0770 0.0537 0.1120 0.0649
+ QueRec CC (0.1) 0.0346 0.0194 0.0542 0.0257 0.0622 0.0381 0.0910 0.0473 0.0680 0.0427 0.0998 0.0530
+ QueRec RRF 0.0338 0.0202 0.0532 0.0264 0.0608 0.0390 0.0920 0.0491 0.0731 0.0483 0.1073 0.0593
ELMRec 0.0492 0.0414 0.0569 0.0437 0.0610 0.0503 0.0729 0.0540 0.0706 0.0616 0.0749 0.0623
+ QueRec CC (Ours) 0.0589 0.0455 0.0746 0.0506 0.0816 0.0625 0.1060 0.0704 0.1092 0.0777 0.1350 0.0860
+ QueRec CC (0.9) 0.0256 0.0170 0.0381 0.0211 0.0477 0.0331 0.0648 0.0386 0.0658 0.0445 0.0919 0.0529
+ QueRec CC (0.7) 0.0337 0.0217 0.0532 0.0279 0.0592 0.0401 0.0867 0.0489 0.0831 0.0564 0.1172 0.0674
+ QueRec CC (0.5) 0.0586 0.0462 0.0717 0.0504 0.0805 0.0617 0.1026 0.0689 0.1032 0.0811 0.1267 0.0887
+ QueRec CC (0.3) 0.0504 0.0427 0.0594 0.0456 0.0665 0.0545 0.0821 0.0595 0.0786 0.0676 0.0936 0.0724
+ QueRec CC (0.1) 0.0491 0.0444 0.0562 0.0421 0.0623 0.0515 0.0745 0.0555 0.0681 0.0621 0.0721 0.0634
+ RRF 0.0576 0.0447 0.0720 0.0494 0.0791 0.0595 0.1014 0.0667 0.0957 0.0752 0.1208 0.0833
Table 6: Experimental results based on various CF-LLM cooperation reranking methods. The best and second-best results for each metric are highlighted in bold and underlined, respectively. (Ours) denotes our proposed Hit Rate-based reranking method, while the other numerical values represent fixed lambda ($\lambda$) values used in the linear combination approach. "H@k" and "N@k" denote Hit Rate and NDCG at rank k, respectively.
To integrate the top-k selection results from both the traditional recommendation model and the LLM-based approach, we employed a linear combination method. To prevent performance imbalance between the two models from negatively affecting the final ranking, we determined the $\lambda$ value based on validation Hit@10 scores and the intersection ratio. To evaluate the effectiveness of this approach, we compared the performance of our method against two alternatives: fixing the $\lambda$ value and employing Reciprocal Rank Fusion (RRF). Table 6 presents the results of these experiments. The results indicate that our proposed method achieved the highest performance in most cases.
Appendix C Prompt and Generated Query
This section presents the prompts used in our query generation approach along with examples of the generated queries. We designed the prompt format based on the LLaMA3 instruction format. For user history, we set the maximum length to 8 and included a phrase to emphasize the last interaction. To prevent excessive prompt length, we incorporated item queries only for the last interaction during user query generation. We instructed the LLMs to generate 10 queries per each item and user. We utilized vLLM  [^4] https://github.com/vllm-project/vllm to perform the generation steps required for training-free methods, including query generation. The complete set of generated queries will be released through an online public repository upon acceptance.
C.1 Item Query Generation Prompt
<|begin_of_text|><|start_header_id|>system<|end_header_id|>
You are an intelligent assistant designed to create detailed and precise search queries for items based on their descriptions and aggregated user reviews.
Your task is to generate 10 distinct and comprehensive search queries that effectively help users find the specified item.
Focus on incorporating key features, standout aspects, brand, and practical benefits into each query to enhance search accuracy.
Emphasize the unique attributes that differentiate the item from similar products.
Each query should be concise, factual, and separated by line breaks.
<|eot_id|>
<|start_header_id|>user<|end_header_id|>
### Task:
Analyze the provided item metadata and user reviews to generate 10 detailed and objective search queries for the item.
Your goal is to create queries that highlight the items key features, benefits, and unique aspects based on its description and user feedback.
### Input:
- **Item Title**: Last Night on Earth: Growing Hunger Expansion
- **Brand**: Flying Frog Productions
- **Description**: The Growing Hunger Expansion introduces new game mechanics and three exciting new Scenarios to challenge players as well as a two player mini-game. Take control of four new Heroes, each with a highly-detailed plastic miniature as well as seven new Red Zombies for use as Plague Carriers, Grave Dead, or to increase the Zombie Horde. New modular game board sections expand the town and feature unique buildings such as the Supermarket, Library, and Antique Shop. New game cards give Zombies a chance to steal weapons from the Heroes and add powerful Double-Handed weapons to the Heroes arsenal, such as Garden Shears and the Fence Post. Also included are two new full color, die-cut counter sheets adding Free Search Markers for the Heroes as well as many more fun ideas to the Last Night on Earth toolbox for limitless use with official web content or creating your own new Scenarios.
-**User Reviews**:
1. Four Stars Great games
2. Last Night on Earth is the best! Last Night on Earth: Growing Hunger is an expansion for the Last Night on Earth game. I love it because it is so much fun! My friends and I get together once a month for game night and this is one we always pull out. This expansion add heroes, props, locations and scenarios to the main game. You do need the main Last Night on Earth board game for this add on to help you in any way. If you would like to know more about the main game there is a video on you tube from Wil Wheaton on his tabletop gaming blog style show called "TableTop" they played last night on earth https://www.youtube.com/watch?v=UhLU2-BuhMIIts a great explanation of the game and how it works! However I took it to the next level and I hand painted my hero figures. I will post the images.
3. Good expansion Its a good expansion for a good game. I dont know what else to write so Im finishing with filler.
4. Great expansion to a great game A worthy expansion to the original game, adding new characters, cards, scenarios, and lots and lots of extra components for building your own scenarios and adventures. For me, thats the big difference between owning just the game and this expansion: whether or not you intend to make your own scenarios or just play the ones that come with the game. If you are one of those "game tinkerers" like me, then you will love this expansion.
5. A good addition to a great game. Ill keep it pretty short. If youre considering the expansion, I would hope you already own Last Night on Earth. I think LNOE is a great game that offers suspense, teamwork, competition, and usually a lot of laughs. This expansion does not change much. What it does do is add some more locations (in the form of some new L-shaped boards), some new characters, new hero and zombie cards, and a few new scenarios. On top of that, it comes with new game pieces that are used with the free scenarios available on the games official website or that can be used to design your own scenarios.In short, this expansion does not reinvent the wheel. What it will do, however, is help expand and reinvigorate an already great game. Its a little pricey, but if you like LNOE, its a good investment.
6. great game add on The is a great game expansion to the game last night on earth, its very fun and different i highly recommend this add on. love the new game pieces
### Requirements:
- Generate exactly 10 distinct search queries, each on a **separate line**.
- Incorporate key metadata and review insights to create effective, descriptive queries.
- Highlight the items purpose, standout features, brand, and practical benefits in each query.
- Emphasize the **unique attributes that differentiate the item from other similar products**.
- Avoid redundant details and ensure each query is unique and precise.
### Response:
<|eot_id|><|start_header_id|>assistant<|end_header_id|>
C.2 Item Query Sample
Toys
1. **"Radica 20Q AI Game for Kids and Adults: Color Options and Educational Value"**
2. **"20Q Toy with Smart Features: How Does it Guess What Youre Thinking?"**
3. **"Best Party Game for Adults and Kids: 20Q Review and Comparison"**
4. **"Portable and Durable 20Q Game for Travel and Classroom Use"**
5. **"Radica 20Q AI Game: How Accurate is it in Guessing Words and Concepts?"**
6. **"Fun and Educational Toy for Kids: 20Q Review and Gift Ideas"**
7. **"20Q Game for Large Groups: Conversation Starter and Icebreaker"**
8. **"Best Gift for Kids and Adults: 20Q Artificial Intelligence Game with Reviews"**
9. **"20Q Toy with Interactive Features: How Does it Engage Kids and Adults?"**
10. **"Radica 20Q Game: Unique Features and Benefits for Home and Classroom Use"**
Beauty
1. **"Fashionwu 10 Pair Long Black False Eyelashes Makeup for Everyday Use, Natural Look with 1cm-1.5cm Length and 3cm Width"**
2. **"Long Black False Eyelashes for Beginners, 10 Pairs with Curved Shape, Perfect for Enhancing Natural Lashes"**
3. **"Fashionwu Black False Eyelashes for Cosplay, 10 Pairs with Thick and Sturdy Design, Great for Drag Queens and Performers"**
4. **"Affordable 10 Pair Long Black False Eyelashes, Good for Special Occasions and Everyday Wear, Natural and Sassy Look"**
5. **"Fashionwu Black False Eyelashes with Glue Strip, 10 Pairs for Long-Lasting Wear, Perfect for Lash Enthusiasts"**
6. **"Long Black False Eyelashes for Beginners, 10 Pairs with Thicker Band and Longer Length, Great for Volume and Drama"**
7. **"Fashionwu Black False Eyelashes for Stage Performers, 10 Pairs with Sturdy Design and Long-Lasting Adhesive"**
8. **"Natural and Sassy Black False Eyelashes, 10 Pairs with Curved Shape and Thicker Band, Perfect for Everyday Wear"**
9. **"Fashionwu Black False Eyelashes for Special Occasions, 10 Pairs with Glitter and Sparkle, Great for Cosplay and Parties"**
10. **"Long Black False Eyelashes for Lash Enthusiasts, 10 Pairs with Thicker Band and Longer Length, Perfect for Volume and Drama"**
Sports
1. **Trijicon NS 3Dot Set (GR/GR) T0963 for Glock: Military-Tested, Bright & Tough Night Sights with 12-Year Tritium Life**
2. **Best Trijicon Night Sights for Glock: 3Dot Set with White Rings for Enhanced Daylight Visibility and Shock Protection**
3. **Trijicon NS 3Dot Set (GR/GR) T0963: Proven Reliability and Accuracy for Combat Handguns**
4. **Trijicon Tritium Night Sights for Glock: 3Dot Set with Aluminum Housing and Silicon Rubber Cushions for Low-Light Performance**
5. **Trijicon NS 3Dot Set (GR/GR) T0963: 12-Year Warranty, Military-Grade Night Sights for Glock and Other Handguns**
6. **Trijicon NS 3Dot Set (GR/GR) T0963: Fast Target Acquisition and Superior Daylight Visibility with White Rings**
7. **Trijicon NS 3Dot Set (GR/GR) T0963 for Glock: Shock-Resistant and Durable Tritium Night Sights**
8. **Trijicon Tritium Night Sights for Glock: 3Dot Set with Enhanced Low-Light Performance and Long-Lasting Tritium Life**
9. **Trijicon NS 3Dot Set (GR/GR) T0963: Fast and Accurate Night Sights for Glock Handguns with White Rings and Aluminum Housing**
10. **Trijicon NS 3Dot Set (GR/GR) T0963: Military-Tested, Proven Night Sights for Glock and Other Handguns with 12-Year Warranty**
C.3 User Query Generation Prompt
C.3.1 Prompt Format
<|begin_of_text|><|start_header_id|>system<|end_header_id|>
You are an intelligent assistant designed to analyze a users purchase history and behavior to generate **10 effective search queries** for predicting the **next items** they are most likely to purchase.
Your task is to evaluate past purchase patterns, item metadata, and related search queries to construct concise and accurate search queries that can be used to find the next recommended items.
Focus on identifying recurring patterns, shifts in preferences, and evolving interests to enhance the relevance of the search queries.
For the most recent item, make sure to include **related queries** that were associated with it to improve search accuracy.
Ensure each query highlights the **unique characteristics of items** and reflects the **users preferences and interests** for more personalized recommendations.
Ensure each query is clear, specific, and optimized for retrieving relevant items.
<|eot_id|>
<|start_header_id|>user<|end_header_id|>
### Task:
You are an intelligent assistant tasked with generating **10 optimized search queries** to predict the **next items** a user is likely to purchase based on their chronological purchase history, item metadata, and related search queries.
**Purchase History:** A chronological list of items the user has purchased, including item brands, categories, descriptions, associated metadata, and related search queries. For the **most recent item**, related queries are also provided to enhance search relevance.
### Output Format:
Your response should follow this exact format, ensuring:
1. Each search query is presented on a **separate line**.
2. **No newlines or additional formatting** within each query.
3. The queries should be concise, specific, and optimized for accurate item retrieval.
### Requirements:
- Generate **10 precise search queries** based on the users purchase history, item metadata, and related search queries.
- For the **most recent item**, ensure that **related queries** are incorporated to improve relevance.
- Ensure each query captures key patterns, preferences, and interests derived from the provided data.
- **Highlight the unique characteristics of items** (e.g., special features, distinctive attributes) and reflect the **users preferences and behavioral trends** in the queries.
- Do **not** include explanations, introductions, or follow-up comments.
- Keep each query **clear, concise, and limited to a single line**.
### Input:
{user_history}
{generated_queries}
### Output:<|eot_id|><|start_header_id|>assistant<|end_header_id|>
C.3.2 User History Example
Toys
**Title:** Melissa &amp; Doug Wooden Take Along Tool Kit (24pc)‘
**Brand:** Melissa &amp; Doug
**Categories:** Toys & Games, Dress Up & Pretend Play, Pretend Play, Construction Tools
**User Review:**
Daughter loves to be like daddy My little girl loves to play pretend and "help" her Daddy and Poppa. She loves to build, hammer, and screw nails and bolts along with them. We have had it about a year and the nails are falling apart - its not quite as well built as most of the Melissa and Doug products we have had in the past. However, we are still very pleased with the product overall. I especially like the fine motor skills it helps to develop.
**Title:** Melissa &amp; Doug Penguin Plush
**Brand:** Melissa &amp; Doug
**Categories:** Toys & Games, Stuffed Animals & Plush, Animals & Figures
**User Review:**
One of her favorite stuffed animals My parents bought this for my daughter last Christmas, and she still loves it. It was as tall as she was last year, so she was adorable at Christmas hugging it and lugging it around the house. Now that she is older, she still loves playing with it and it is easier for her to manipulate.One reason they bought the penguin was due to her love of the movie "Happy Feet". One of her favorite things to do is lie on the floor with her penguin while watching the movie. I dont see much educational value other than supporting (or spurring) an interest in penguins.
**Title:** Melissa &amp; Doug Farm Sound Puzzle
**Brand:** Melissa &amp; Doug
**Categories:** Toys & Games, Puzzles, Pegged Puzzles
**User Review:**
Another favorite This is another great Melissa and Doug product my parents bought for my daughter. When she was two, she got this for her birthday. It was so cute to see her face the first time she lifted one of the pieces and heard the animal sound. Personally, we thought the sound quality was fine.The wooden construction provides plenty of durability, and she did not tear it up. We were able to pass it on to her younger cousin. The sound quality was deteriorating slightly by that time, but it was still relatively easy to hear. We had it for about one year before passing it to her cousin.
This is the most recently purchased product:
**Title:** Melissa &amp; Doug Flip to Win Memory Game
**Brand:** Melissa &amp; Doug
**Categories:** Toys & Games, Games, Board Games
**User Review:**
Another great product I bought this last year for my now four-year-old daughter. We have always been pleased with the Melissa and Doug product line, and this one does not disappoint. I thought the cardboard "sheets" that you interchange would not last, but they have held up surprisingly well. We only have one that is a bit dog-eared, and thats due to my daughter thinking it would taste delicious so into the mouth it went.Its been a great learning tool. For example, the colors let us explore colors far beyond the basic 6 or 8 most learning tools address. This has increased her vocabulary, and it has made learning much more fun.All in all, this is one of my favorite products - and my daughter likes it a ...
Beauty
**Title:** Now Foods: Tea Tree Oil, 4 oz
**Brand:** Now Foods
**Categories:** Beauty, Skin Care, Body, Moisturizers, Oils
**User Review:**
For keloid scars. I bought this to treat my keloid scars and well, it didnt work, but I give it a high rating anyway because it does have other benefits and uses. The smell isnt as bad, it smells like wet wood or wet bark.
**Title:** Dermablend Quick Fix Concealer SPF 30, Ivory
**Brand:** Dermablend
**Categories:** Beauty, Makeup, Face, Concealers & Neutralizers
**User Review:**
Works well but wrong color. The color described isnt accurate. This turned out to be a more pinkish color and I am tan/yellowish so it did not work well for me. It is really thick though so it works as far as covering blemishes.
**Title:** Finulite Cellulite Smoothing Massage Mitt
**Brand:** Finulite
**Categories:** Beauty, Bath & Body, Bathing Accessories, Bath Mitts & Cloths
**User Review:**
Smooth skin. I used this on my thighs with the Finulite cream and in the shower with regular body soap and it made my skin super smooth. It removes dead skin and you can see it on the massage mitt. Its easy to clean as well. Its a tad painful on dry skin but in the shower when your skin is wet, it glides very smoothly and feels really nice.
**Title:** Xtreme Brite Brightening Gel 1oz.‘
**Brand:** Xtreme Brite
**Categories:** Beauty, Hair Care, Styling Products, Creams, Gels & Lotions
**User Review:**
Mixed feelings. I have mixed feelings about this product. When I first started using it, I can see and feel the difference. It actually started to literally peel the dark skin away. Theres a bit of a sting but thats how I know it was working. My skin was lighter in just a couple weeks! However, once Ive stopped using it when I reached my desired skin color, it went back to being dark. I applied more and suddenly it was no longer working. It didnt peel or sting. I wonder what went wrong or if it has an expiration date or something. Overall, it works but definitely not consistent.
**Title:** Remington CI95AC/2 Tstudio Salon Collection Pearl Digital Ceramic Curling Wand, 1/2 Inch - 1 Inch
**Brand:** Remington
**Categories:** Beauty, Hair Care, Styling Tools, Irons, Curling Irons
**User Review:**
My favorite tool! These are great for natural, beachy waves for your hair. The heatproof gloves work wonders as well. It heats up real hot and is super easy to use.
This is the most recently purchased product:
**Title:** Finulite - The End to Cellulite, AM/PM Cellulite Cream (2 - 4 oz tubes)‘
**Brand:** Unknown
**Categories:** Beauty, Skin Care, Body, Moisturizers, Creams
**User Review:**
I guess its my fault. The main thing with these creams is you have to be diligent and consistent, which I wasnt. It is a PAIN in the butt to do it every night and every morning, scrubbing with the somewhat painful glove I bought with it. It did make my skin super smooth, but cellulite is a bitch to get rid of even with exercise (lots of thin people I know have cellulite, so its definitely not a fat people thing). Long story short, I got tired of the routine and gave up.
Sports
**Title:** Tone Fitness Cement Filled Kettlebell Set - 30 lbs.‘
**Brand:** Tone Fitness
**Categories:** Sports & Outdoors, Exercise & Fitness, Strength Training Equipment, Kettlebells
**User Review:**
EXCELLENT Prouct EXCELLENT kettlebells for both men and women. Vinyl is very smooth and easy on the hands. I would highly recommend.
**Title:** Blackburn AirTower 2 Bicycle Pump, Silver
**Brand:** Unknown
**Categories:** Sports & Outdoors, Cycling, Accessories, Bike Pumps, Floor Pumps
**User Review:**
Dy-no-mite This think can fill an auto tire in minutes. The best pump I have ever had. Highly recommend it. Not cheap thin metal. This is good quality thick steel.
**Title:** Kimber Pepperblaster 2 Red, One Size
**Brand:** Kimber
**Categories:** Sports & Outdoors, Outdoor Gear, Camping & Hiking, Personal Care
**User Review:**
Welll now... This shoots two shots of Pepper spray. That is it. You throw it away. It is NOT reloadable. It is the same excellent Kimber quality as their weapons are... But how do I know where to aim if I cant practice shoot it ??? I would not recommend this to anyone.
**Title:** Ultimate Arms Gear Tactical 4 Reticle Red Dot Open Reflex Sight with Weaver-Picatinny Rail Mount
**Brand:** Ultimate Arms Gear
**Categories:** Sports & Outdoors, Hunting & Fishing, Hunting, Hunting Optics, Gun Scopes, Rifle Scopes
**User Review:**
Excellent Fits my gun great. Easy to mount. Very large so the image is clear. No reason you should ever miss a target with this site. GREAT quality.
**Title:** Camelbak Podium Big Chill 25 oz Bottle
**Brand:** CamelBak
**Categories:** Sports & Outdoors, Accessories, Sports Water Bottles
**User Review:**
Ahhh.... Hard to get enough water out of this bottle. Doesnt keep anything cold. Very thin material. I am worried after being out in the sun, how long before it breaks.
**Title:** Kimber Pepperblaster Ii Holster
**Brand:** Meprolight
**Categories:** Sports & Outdoors, Paintball & Airsoft, Airsoft, Holsters
**User Review:**
As I said in the other feedback If you think you need the Kimber Pepperblaster then this is a great holster to carry it in. Now that I found out you cant reload you Pepperblaster, I see no reason for either the blaster or the holster. How do you practice ...???
**Title:** ‘5 LED Bicycle Rear Tail Red Bike Torch Laser Beam Lamp Light
**Brand:**
**Categories:** Sports & Outdoors, Cycling, Lights & Reflectors, Taillights
**User Review:**
Great lite Very bright. I would recommend this lite to anyone. I put one on my wifes bike. I love this light.
This is the most recently purchased product:
**Title:** Body Solid MA307N Nylon Head Harness
**Brand:** Body Solid
**Categories:** Sports & Outdoors, Exercise & Fitness, Exercise Machine Accessories, Exercise Machine Attachments
**User Review:**
Exactly what I wanted My friend runs a boxing school for urban kids. These kids work hard at staying out of trouble, doing their school work and being able to learn to box. This harness works perfect. It is of highest quality.
C.4 User Query Sample
C.4.1 Toys
1. "LeapFrog TAG Reading System Purple 32MB pen with phonetic sound and games for 4-year-olds, educational toy for learning to read and spell"
2. "Best LeapFrog reading system for preschoolers with built-in games and activities, suitable for travel and home use"
3. "LeapFrog TAG Reading System Purple review, educational toy for kids to learn alphabet, spelling, and reading skills with phonetic sound and games"
4. "LeapFrog TAG Reading System 32MB pen with LeapFrog Connect software for downloading and managing content, suitable for children aged 3-6 years"
5. "LeapFrog TAG Reading System Purple with headphones, compact and portable learning tool for kids to learn to read and spell with fun games and activities"
6. "LeapFrog TAG Reading System Purple review, educational toy for kids to learn phonics, sight words, and reading skills with LeapFrog Connect software and games"
7. "LeapFrog TAG Reading System Purple 32MB pen with built-in games and activities, suitable for kids to learn to read and spell with phonetic sound and fun stories"
8. "LeapFrog TAG Reading System Purple with LeapFrog Connect software for downloading and managing content, suitable for children aged 3-6 years, educational toy for learning to read and spell"
9. "LeapFrog TAG Reading System Purple review, educational toy for kids to learn alphabet, spelling, and reading skills with phonetic sound, games, and activities"
10. "LeapFrog TAG Reading System Purple 32MB pen with LeapFrog Connect software and games, suitable for kids to learn to read and spell with fun stories and activities, compact and portable"
C.4.2 Beauty
1. "Neutrogena Microdermabrasion System with glycerin and ultra-fine crystals for smooth, luminous skin and anti-aging benefits, clinically proven for visible results in one use."
2. "Best affordable microdermabrasion system for acne-prone skin, gentle exfoliation, and firming, with 12 pre-dosed puffs and AA batteries included."
3. "Neutrogena Microdermabrasion System with massaging micro-vibrations for improved skin texture, radiance, and fine line reduction, suitable for sensitive skin types."
4. "Microdermabrasion system for at-home use, with dermatologist-recommended Neutrogena brand, proven to deliver smoother, more luminous skin in just one use, and affordable price point."
5. "Exfoliating microdermabrasion system with gentle, pre-dosed puffs and soothing glycerin, ideal for daily use, with visible results in just one treatment, and suitable for all skin types."
6. "Neutrogena Microdermabrasion System with anti-aging benefits, firming, and skin brightening, with a unique combination of exfoliating crystals and micro-vibrations, and easy to use at home."
7. "Best microdermabrasion system for oily skin, with a gentle, non-irritating formula, and visible results in just one use, with a affordable price and convenient packaging."
8. "Neutrogena Microdermabrasion System with a dermatologist-recommended formula, proven to deliver smoother, more radiant skin, and suitable for daily use, with a unique combination of exfoliation and micro-vibrations."
9. "Microdermabrasion system for anti-aging, firming, and skin brightening, with a gentle, pre-dosed puff system, and visible results in just one treatment, and suitable for sensitive skin types."
10. "Neutrogena Microdermabrasion System with a clinically proven formula, delivering visible results in just one use, and a affordable price point, with a unique combination of exfoliation and micro-vibrations for smoother, more luminous skin."
C.4.3 Sports
1. "ProSource Heavy-Duty Easy Gym Doorway Chin-Up/Pull-Up Bar with 300lb weight capacity and multi-position design for home workout"
2. "Best doorway pull-up bar for thin walls and easy installation with ProSource Comfort Grip technology"
3. "Heavy-duty pull-up bar for home gym with adjustable grip and sturdy construction for 24-32 inch doorways"
4. "ProSource Easy Gym Doorway Chin-Up/Pull-Up Bar with wall-mounting option and 5-star durability rating"
5. "Inexpensive and easy-to-assemble pull-up bar for home workout with 300lb weight capacity and multi-functional design"
6. "Best pull-up bar for doorways with raised molding and minimal wall damage with ProSource brand guarantee"
7. "Heavy-duty doorway pull-up bar with adjustable grip and comfortable grip technology for home gym workouts"
8. "ProSource Heavy-Duty Easy Gym Doorway Chin-Up/Pull-Up Bar with 300lb weight capacity and easy installation for home fitness"
9. "Best pull-up bar for doorways with multi-position design and sturdy construction for home workout and exercise"
10. "ProSource Easy Gym Doorway Chin-Up/Pull-Up Bar with 300lb weight capacity and wall-mounting option for home gym and fitness enthusiasts"